# Can LLM-Simulated Practice and Feedback Upskill Human Counselors? A Randomized Study with 90+ Novice Counselors

### v1

@@ 4.5.1 Self-Efficacy Measurement Instruments
The final instrument comprised 12 items across three factors: (1) exploration skills (e.g., restatements, reflecting feelings, open questions, listening); (2) actions skills (e.g., providing suggestions, knowing which actions to take); and (3) session management skills (e.g., keeping sessions on track). Reliability analysis demonstrated good to excellent internal consistency across all factors, with Cronbach’s $\alpha$ values of 0.784, 0.803, and 0.905 for ES, AS, and SM, respectively.
4.5.2 Dunning-Kruger Effect Analysis
Our primary analysis investigated potential mis-calibration between participants’ self-assessments and their actual counseling performance, specifically examining whether data exhibited patterns consistent with the Dunning-Kruger effect. This phenomenon [38] suggests that individuals with lower skill levels tend to overestimate their abilities, while highly skilled individuals may slightly underestimate their competence. We focused this analysis on Exploration Skills (ES) and Action Skills (AS), as these dimensions had straightforward mappings between CASES items and our NLP behavioral classifiers (Table 3).
To test for the presence of the Dunning-Kruger effect, we follow the classic analysis method that splits the data into quartiles based on performance and conducts a two-way analysis of variances for self-assessments and actual performance and across the quartiles; and finally verifies via post-hoc tests that the bottom performers have the biggest overestimation of their abilities [38]. To standardize the comparison between self-efficacy and performance, we transform each measure into a percentile rank (0 - 99) computed across all data collected for the pre-intervention and post-intervention chats ($b_{0},b_{1}$; $s_{0},s_{1}$).
4.5.3 Calibration Change Analysis
To evaluate whether our interventions improved self-efficacy calibration, we computed discrepancy scores by subtracting standardized performance scores from standardized self-efficacy scores for each participant at both assessment timepoints. These discrepancy metrics provided a direct measure of calibration, with positive values indicating overconfidence and negative values indicating under-confidence.
We then examined changes in these discrepancy scores from pre- to post-intervention for both intervention groups. This analysis allowed us to determine whether either intervention significantly improved the alignment between participants’ self-perceptions and their actual counseling abilities.
4.5.4 Changes in Raw Self-Efficacy

### v2

@@ 4.4. Measures
The final instrument comprised 12 items across three factors: (1) exploration skills (e.g., restatements, reflecting feelings, open questions, listening); (2) action skills (e.g., providing suggestions, knowing which actions to take); and (3) session management skills (e.g., keeping sessions on track). To assess the internal consistency of each factor, we conducted reliability analysis using Cronbach’s $\alpha$, which measures how closely related a set of items are as a group (Tavakol and Dennick, 2011). The analysis demonstrated good to excellent internal consistency across all factors, with Cronbach’s $\alpha$ values of 0.784, 0.803, and 0.905 for exploration skills, action skills, and session management skills, respectively.
Stages Skill Category Description
Table 2. Overview of our analysis of skill development, grounded in Hill’s Helping Skills model (Nelson-Jones, 2013). We select a skill subset relevant for beginning counselors at the undergraduate and first-year graduate level (Hunsmann et al., 2024). These include microskills during the exploration and action stages; and macro skills that are applicable throughout the session. Hill’s insight stage, of which self-disclosure was the only relevant skill for basic counseling, was excluded from our primary analyses due to its infrequent occurrence in our data.
4.4.3. RQ3. Qualitative Self-Reflections on Therapeutic Intentions
LLM-simulated training provides opportunities for experiential learning (Huerta-Wong and Schoech, 2010) whereby reflection on action (Schön, 2017) can support counselors in refining their therapeutic intentions and strategies. To study this impact on participants’ intentions, we collected qualitative self-reflections from two time points: immediately after the training intervention chat, where participants responded to ”What would you do differently as a therapist?” and after the post-intervention chat, where they reflected on ”What did you do well as a therapist?”. We examined how initial intentions translated into reported strengths across the P+F and P groups.
4.4.4. RQ4. Quantitative Experience of CARE’s LLM Feedback and Simulated Practice
Three survey questions measured participants’ perceptions of CARE’s AI feedback system on a 5-point Likert scale. Helpfulness. Participants rated ”To what extent do you find the AI feedback to be constructive and helpful?”. Comfort. Participants rated ”To what extent do you agree with the following statement: ’I am comfortable receiving AI feedback’”. Readiness. Participants rated ”The AI feedback system is ready to be used by counselors-in-training.” Participants in the P+F group answered these questions after the intervention chat. Participants in the P group also received feedback–only after the pre-post experimental measures were completed–and subsequently answered these three questions about AI feedback.
We measured participants’ perceptions of each of the AI patients after each chat (pre-intervention, practice intervention, post-intervention) with several 7-point Likert scale items. Authenticity. Participants rated ”The AI patient was authentic in its role.” Four questions from the NASA-TLX workload scale were given after each simulated practice: Mental Demand: ”How mentally demanding was giving counseling support to this patient?”; Temporal Demand: ”How hurried or rushed did you feel giving counseling support to this patient?”; Effort: ”How hard did you have to work to accomplish your level of performance”; Frustration: ”How discouraged or stressed were you while giving counseling support to this patient?”.
4.4.5. RQ5 and RQ6. Qualitative Experience of using CARE’s LLM Feedback and Simulated Practice
Participant interview data was collected at different time-points throughout the 75-minute session. To specifically understand the experience receiving feedback from CARE (RQ5), participants were asked to elaborate in more detail their answers after filling out the three survey questions about CARE’s feedback; re-review the AI feedback page and think-aloud about their agreements or disagreements with any of the feedback; and explain whether the feedback had any bearing on their self-reflections about what they did well or wanted to do differently as a counselor. To specifically understand the experience practicing with CARE’s simulated patients (RQ6), participants were given the chance to explain in more detail their answers to the quantitative survey items about the simulated patients. Finally, a semi-structured exit-interview was conducted for all participants which was framed around the following questions: ”What do you like about this training tool for helping skills?” ”What do you wish was different about the training tool?” and ”What suggestions do you have for improving any part of the training tool?”.
4.5. Analyses
4.5.1. RQ1. Effects on Behavioral Performance
Our analysis of behavioral performance consists of two perspectives: (1) testing changes in behaviors of skills used across time (pre-intervention vs. post-intervention) and between intervention groups (P vs. PF); and (2) analyzing the relationship between intervention-exposure to good alternative patterns in AI feedback and post-intervention skill use.
Testing Changes Across Time and Between Groups
Using our selected classifiers, we examined skill usage changes from pre- to post-intervention. For each transcript, we computed the proportion of utterances showing strong skill use ($b=U_{\text{strengths}}/U_{\text{total}}$) or needing improvement ($b=U_{\text{improvement}}/U_{\text{total}}$).
To test for pre-post changes ($b_{0},b_{1}$), we used paired $t$-tests and Cohen’s $d$ effect sizes. To compare P and P+F groups ($b_{1}^{P}-b_{0}^{P}$ vs. $b_{1}^{PF}-b_{0}^{PF}$), we used unpaired $t$-tests. We conducted 12 planned $t$-tests: three skills (Empathy, Reflections, Questions) for strong uses and one skill (Suggestions) for areas needing improvement, analyzing both within-group changes and between-group differences.
Exposure to Good Alternatives in AI Feedback
To assess whether AI feedback exposure affected post-intervention performance, we defined Good Alternatives during Practice (GAP) as the proportion of trainee utterances for which the AI suggested an alternative response that exemplified a strong use of a skill:
$T_{GAP}=\frac{A_{strengths}}{U_{total}},$
where $A_{strengths}$ is the number of AI-generated alternative responses judged as strong exemplars and $U_{total}$ is the total trainee utterances. We then fit a lagged linear regression predicting post-chat behavior $b_{1}$ while controlling for pre-chat behavior $b_{0}$ and including GAP as a predictor:
$b_{1}=\beta_{0}+\beta_{1}b_{0}+\beta_{2}T_{GAP},$
thereby isolating the effect of feedback exposure.
4.5.2. RQ2. Effects on Self-Efficacy and its (mis)calibration with Behavioral Performance
First, we test for changes in raw self-efficacy scores after practice or practice-and-feedback. Second, we examine the calibration of self-efficacy ratings with actual performance. Third, we evaluate whether P or P+F interventions improve this calibration.
Changes in Raw Self-Efficacy